# Semiclassical descriptions of dissipative dynamics of strongly interacting Bose gases in optical lattices

### v2

@@ IV.1 Stability of the unit filling Mott insulating state
Figure 2: (a) SU(3) dTWA results for the non-unitary time evolution of the atomic density $n_{j}(t)$ starting from the unit filling Mott insulating state $|\psi_{\rm Mott}\rangle$ at time $t=0$ for three different values of the dissipation strength $\gamma=\hbar\Gamma_{\rm PA}/U$. The results for $\gamma=0.1$, $1$, and $5$ are shown by the blue solid, red dotted, black dashed lines, respectively. (b) Same as (a) but for the SU(3) FP-TWA results. (c) Two-body loss rate $\kappa$ as a function of $\gamma$ (blue open squares) extracted from the dTWA results for $n_{j}(t)$ shown in (a). The error bar for each result denotes the standard estimation error for the fitting. We select the data points of density $n_{j}(t)$ in (a) for the fitting such that the distance between adjacent data points on the time axis is $\delta t_{\rm fit}=2.0\hbar/U$. The black solid line is the mean-field result $\kappa_{\rm MF}$ of a dissipative hard-core boson model. (d) Same as (c) but for the SU(3) FP-TWA results. For comparison, the experimental results reported in Ref. Tomita et al. (2017) are also plotted as red open circles in (c) and (d). Here, in all calculations, we set the hopping amplitude $J=0.08U$.
@@
In order to extract the two-body loss rate $\kappa$ from the behavior of the decay in $n_{j}(t)$, we utilize a fitting function of the form $f_{\kappa}(t)=(1+\kappa t)^{-1}$ for the early time regimes satisfying $n_{j}(t)>0.4$ Tomita et al. (2017). Note that this fitting function is the solution of a damping equation of the density associated with two-body losses, i.e., $\frac{dn_{j}}{dt}=-\kappa n^{2}_{j}$ with the initial conditions $n_{j}(t=0)=1$ Syassen et al. (2008). Figure 2(c) shows the loss rate $\kappa$ for the SU(3) dTWA as a function of $\gamma$. Note that $\gamma$ is independent of the optical lattice depth. Here we assume the recoil energy of the optical lattice $E_{\rm R}/\hbar\approx 2\pi\times 4.0$ kHz for the lattice constant $d_{\rm lat}=266$ nm. One of the main findings in Fig. 2(c) is the manifestation of the continuous quantum Zeno effect in the vicinity of $\gamma\sim 1$, which is qualitatively consistent with the experimental observation in Ref. Tomita et al. (2017). For more quantitative comparison, we also show the corresponding experimental results in the same figure and find that they are in good agreement with our theoretical results within the experimental error bars. In the weak dissipation regime with $\gamma\lesssim 1$, the loss rate $\kappa$ becomes larger with increasing $\gamma$. In constract, in the strong dissipation regime with $\gamma\gtrsim 1$, the loss rate $\kappa$ decreases with increasing $\gamma$, which can be attributed to the suppression of doubly occupied sites due to strong two-body losses, i.e., the Zeno effect Tomita et al. (2017).
@@
Moreover, the loss rate estimated within the SU(3) dTWA is much better than the mean field result $\hbar\kappa_{\rm MF}=16z\gamma J^{2}U^{-1}/(4+\gamma^{2})\equiv 4z\hbar\Gamma_{\rm eff}$ from a previous work García-Ripoll et al. (2009), where $z=2d=6$ is the coordination number of the lattice geometry. This mean field result is obtained by performing a site-decoupling approximation for the dynamical equation of $\langle{\hat{n}}_{j}\rangle$ with mapping to a hard-core boson model from the master equation in Eq. (1). The derivation of this effective model is based on a second-order expansion of the hopping term to adiabatically integrate Fock states, including doubly occupied sites, which is validated under the conditions of strong dissipation and strong interaction, i.e., $J\ll\hbar\Gamma_{\rm PA}$ and $J\ll U$. As shown in Fig. 2(c), the mean field treatment also captures the Zeno effect, and the curve of $\kappa_{\rm MF}$ exhibits a maximum at $\gamma=2$, which is comparable to the dTWA result as well as the experimental result. In particular, the dTWA result and the mean field result are in good agreement for $\gamma\lesssim 1$. However, the mean field result overestimates the loss rate for $\gamma\gg 1$. The significant deviations in the strong dissipative regime imply that the mean field approximation for the hard-core boson representation ignores important effects that suppress loss events more effectively, which are successfully captured in the dTWA.
@@
Figure 3: Rescaled time evolution of the atomic density in units of the effective frequency $\Gamma_{\rm eff}$. (a) The SU(3) dTWA results for $\gamma=1$. The blue solid, red dotted, and black dashed lines correspond to $J/U=0.04$, $0.08$, and $0.1$, respectively. Each line is calculated until $t=100\hbar/U$. For comparison, the mean field result $n_{\rm MF}(t)=1/(1+4z\Gamma_{\rm eff}t)$ is also plotted by a black solid line. (b) Same as (a) but for $\gamma=5$. (c) Same as (a) but for the SU(3) FP-TWA results. (d) Same as (c) but for $\gamma=5$.
@@
Let us further discuss the qualitative difference between these two semiclassical descriptions by comparing the results in more detail. The onsite linearity of the SU(3) dTWA leads to long-time behavior for $tU/\hbar\gtrsim 10$ that is qualitatively distinguishable from that in the SU(3) FP-TWA. In particular, for the typical values of $\gamma$ in Fig. 2, the density $n_{j}(t)$ at, for example, $tU/\hbar=100$ in the SU(3) dTWA significantly deviates from that in the SU(3) FP-TWA, even though the short-time behavior agrees well for $tU/\hbar\lesssim 10$. Moreover, as shown in Fig. 2(b), $n_{j}(t)$ obtained by the SU(3) FP-TWA for $\gamma=1$ and $5$ crosses at a certain time, while there is no such intersection found for the SU(3) dTWA results in Fig. 2(a). To gain insight into the validity of the TWA descriptions, we will compare the semiclassical results for a small cluster with the numerically exact results of the master equation in Sec. IV.4. There, we find that the long-time behavior obtained by the SU(3) dTWA is qualitatively similar to the exact results, and that the SU(3) FP-TWA is less qualitative than the SU(3) dTWA due to the nonlinearity of the dissipative contribution. Therefore, we conclude that the SU(3) dTWA results in Fig. 2 are more reliable.
@@ IV.2 Linear sweep time sequence of the hopping
Figure 4: (a-b) SU(3) dTWA results of the non-unitary time evolution of the atomic density $n_{j}(t)$ for the atom loss dynamics, started from the unit filling Mott insulating state $|\psi_{\rm Mott}\rangle$ at time $t=0$, with the linear sweep time sequence crossing the ground state phase transition point of the Bose-Hubbard Hamiltonian, i.e., $J_{\rm ini}=0$ and $J_{\rm fin}=0.04U$. The values of the dissipation strength $\gamma$ are indicated in the figures. The time grid size for the Runge-Kutta scheme is set to be $\Delta t=2.0\times 10^{-3}\hbar/U$. (c-d) Same as (a-b) but for SU(3) FP-TWA results. Here, the time grid size for the Euler-Maruyama scheme is set to be $\Delta t=2.0\times 10^{-4}\hbar/U$. For comparison, the experimental results in Ref. Tomita et al. (2017) are also plotted by circles with error bars in (a)–(d). (e-f) Time evolution of the phase coherence extracted from the single particle correlation functions, i.e., $S({\bm{0}},t)$, for the linear sweep time sequence evaluated by (e) the SU(3) dTWA and (f) the SU(3) FP-TWA.
@@
Next, we analyze how the two-body loss dissipation affects the coherent time evolution caused by a slow parameter change in the Hamiltonian. At zero temperature and for a commensurate mean filling, the Bose-Hubbard Hamiltonian $\hat{H}_{\rm BH}$ exhibits a second order quantum phase transition between a Mott insulating state and a superfluid state Fisher et al. (1989); Capogrosso-Sansone et al. (2007). In the three dimensional setup with the mean occupation tuned to unit filling, the phase transition occurs at $J_{\rm c}=0.0341U$. This exact critical value was numerically estimated in Ref. Capogrosso-Sansone et al. (2007) by employing a quantum Monte Carlo method. Note that the mean-field Gutzwiller approximation yields the critical point $J_{\rm c}^{\rm MFA}=0.0286U$ van Oosten et al. (2001), which is smaller than the exact value due to the less accuracy of the site-decoupling approximation.
@@
Here, in our simulations, we prepare the unit filling Mott insulating state $|\psi_{\rm Mott}\rangle$ at $t=0$ and then slowly drive it with the Lindblad equation for a time-dependent hopping amplitude at a fixed value of the dimensionless dissipation parameter $\gamma$. To realize a setup that is close to the real experiment, we assume a linear sweep time sequence of the hopping given by $J(t)=J_{\rm ini}+(J_{\rm fin}-J_{\rm ini})(t/\tau)$ for $0\leq t\leq\tau$ Tomita et al. (2017). Here, $J_{\rm ini}$ and $J_{\rm fin}$ are the initial and final hopping amplitudes, which are set in the Mott insulating and superfluid regimes, respectively. During this sweep, the onsite interaction strength $U$ remains constant. In the experiment reported in Ref. Tomita et al. (2017), the tuned parameter is the optical lattice depth, which is linearly ramped down from the Mott insulating regime to the superfluid regime. The dominant effect arising in the dynamics during the ramping down is the exponential increase of the hopping energy. Hence, the linear sweep time sequence of the hopping provides a reasonable simplification of the time sequence realized in the experiment.
@@
Figures 4(a) and 4(b) show the SU(3) dTWA results of the atomic density $n_{j}(t)$ for the atom loss dynamics with the time sequence of the hopping amplitude $J(t)$ described above. We specifically choose $J_{\rm ini}=0$ and $J_{\rm fin}=0.04U$ such that $J_{\rm ini}<J_{\rm c}^{\rm MFA}<J_{\rm c}<J_{\rm fin}$ is satisfied. The sweep time is taken to be $\tau=100\hbar/U$, for which non-adiabatic excitations of the system are sufficiently suppressed during the sweep. For the weak dissipation strength such as $\gamma=0.5$, the loss of atoms becomes significant at an onset time $tU/\hbar\sim 83.5$. Here, the onset time is defined as the time when the atomic density drops below $0.8$, as indicated by a dotted line in Fig. 4(a). As shown in Fig. 4(a), the onset time shifts to an earlier time with increasing $\gamma$ for $\gamma\lesssim 2$. Indeed, the onset times for $\gamma=1$ and $\gamma=2$ are $tU/\hbar\sim 73.9$ and $tU/\hbar\sim 71.1$, respectively. However, for the strong dissipation strength, i.e., $\gamma\gtrsim 2$, the onset time shifts to a later time with increasing $\gamma$, as shown in Fig. 4(b). This result manifests the continuous quantum Zeno effect in the sweep dynamics, and a similar behavior has also been observed experimentally in Ref. Tomita et al. (2017). The threshold strength of $\gamma$ separating the Zeno and non-Zeno regimes is $\gamma\sim 2$, which also agrees well quantitatively with the experimental observation Tomita et al. (2017). Since the numerically exact result for a small cluster also supports a similar onset value, as demonstrated in Sec. IV.4, our result here is considered to be qualitatively reliable.
@@
The SU(3) FP-TWA also describes the delay of the atomic loss in the sweep dynamics for the strong dissipation strength. However, it gives rise to a qualitatively different result, as shown in Figs. 4(c) and 4(d). Namely, in contrast to the dTWA result, the delay occurs already for $\gamma\gtrsim 0.1$, implying that the Zeno regime is broadened on the side of smaller $\gamma$. For instance, along the time slice at $t=40\hbar/U$ in Figs. 4(c) and 4(d), the atomic density is found to be significantly suppressed even for $\gamma=0.1$, and it starts to increase with increasing $\gamma$ beyond the threshold. In other words, the nonlinearity of the FP-type approximation manifests itself in an excessive suppression of the atomic density, thus broadening the Zeno regime in the SU(3) FP-TWA.
@@
In Figs. 4(a)–4(d), we also plot the experimental results of the atomic density for $\gamma=0.5$, $1.1$, $2.3$, and $4.2$ reported in Ref. Tomita et al. (2017), which should be directly compared with our TWA results for $\gamma=0.5$, $1$, $2$, and $4$, respectively. Here, the experimental results are plotted along the time axis such that the values of $J/U$ during the sweep in the experiment are the same as the corresponding values in the linear sweep time sequence employed in our simulations. As shown in these figures, we find that, specifically in the SU(3) dTWA, the qualitative behavior of the $\gamma$ dependence of the melting dynamics is reasonably captured within the experimental error bars.
@@
We further analyze the suppression of the superfluid phase coherence in the presence of the strong two-body loss dissipation. For this purpose, we consider a spatial Fourier transform of the equal-time single particle correlation function, i.e., the momentum distribution, defined by
@@
where ${\bf R}_{i}$ represents the real space vector specifying site $i$ and ${\bf k}$ denotes the momentum vector in three dimensions. Note that this distribution is measurable in experiments by using the time-of-flight (TOF) imaging techniques Lewenstein et al. (2012). Here, we specifically set ${\bf k}=(0,0,0)$ for the purpose of quantifying the emergence of coherence over long distances after the hopping sweep. In addition, we note that no inhomogeneity effect due to a trapping potential for the gas is taken into account here.
Figure 4(e) shows the SU(3) dTWA results for $S({\bf 0},t)$. In the absence of dissipation, i.e., $\gamma=0$, the growth of long distance correlations is recovered within this approximation, implying a dynamical phase transition of the system from the Mott insulating phase to a superfluid phase. For the weak dissipation strength, i.e., $\gamma\leq 1$, this semiclassical method remains to produce large occupation of the zero momentum distribution at the final state, i.e., $L^{-3}S({\bf 0},t=\tau)\gg 1$. Hence, the phase coherence comprised of coherent bosons survives well against the weak dissipation. However, for the strong dissipation strength, i.e., $\gamma\gtrsim 2$, the distribution is found to be strongly suppressed such that $L^{-3}S({\bf 0},t=\tau)\sim 1$, implying that the strong dissipation stabilizes the Mott insulating state and leads to the delay of the evolution toward a superfluid phase due to the continuous Zeno effect. This result is consistent with the experimental observation in TOF absorption images Tomita et al. (2017), revealing that the phase coherence associated with superfluidity is suppressed by the strong dissipation.

### v3

@@ IV.1 Stability of the unit filling Mott insulating state
Figure 2: (a) SU(3) dTWA results for the nonunitary time evolution of the atomic density $n_{j}(t)$ starting from the unit filling Mott-insulating state $|\psi_{\rm Mott}\rangle$ at time $t=0$ for three different values of the dissipation strength $\gamma=\hbar\Gamma_{\rm PA}/U$. The results for $\gamma=0.1$, $1$, and $5$ are shown by bold solid, dotted, and dashed lines, respectively. The corresponding mean-field results with the form $(1+\kappa_{\rm MF}t)^{-1}$ are also plotted as thin black solid, dotted, and dashed lines, respectively. (b) Same as (a) but for the SU(3) FP-TWA results. (c) Two-body loss rate $\kappa$ as a function of $\gamma$ (blue open squares) extracted from the dTWA results for $n_{j}(t)$ shown in (a). The error bar for each result denotes the standard estimation error for the fitting. We select the data points of density $n_{j}(t)$ in (a) for the fitting such that the distance between adjacent data points on the time axis is $\delta t_{\rm fit}=2.0\hbar/U$. The black solid line is the mean-field result $\kappa_{\rm MF}$ of a dissipative hard-core boson model. (d) Same as (c) but for the SU(3) FP-TWA results. For comparison, the experimental results reported in Ref. Tomita et al. (2017) are also plotted as red open circles in (c) and (d). Here, in all calculations, we set the hopping amplitude $J=0.08U$.
@@
In order to extract the two-body loss rate $\kappa$ from the behavior of the decay in $n_{j}(t)$, we utilize a fitting function of the form $f_{\kappa}(t)=(1+\kappa t)^{-1}$ for the early time regimes satisfying $n_{j}(t)>0.4$ Tomita et al. (2017). Note that this fitting function is the solution of a damping equation of the density associated with two-body losses, i.e., $\frac{dn_{j}}{dt}=-\kappa n^{2}_{j}$ with the initial conditions $n_{j}(t=0)=1$ Syassen et al. (2008). Figure 2(c) shows the loss rate $\kappa$ for the SU(3) dTWA as a function of $\gamma$. Note that $\gamma$ is independent of the optical lattice depth. Here we assume the recoil energy of the optical lattice $E_{\rm R}/\hbar\approx 2\pi\times 4.0$ kHz for the lattice constant $d_{\rm lat}=266$ nm. One of the main findings in Fig. 2(c) is the manifestation of the continuous quantum Zeno effect in the vicinity of $\gamma\sim 1$, which is qualitatively consistent with the experimental observation in Ref. Tomita et al. (2017). For more quantitative comparison, we also show the corresponding experimental results in the same figure and find that they are in good agreement with our theoretical results within the experimental error bars. In the weak dissipation regime with $\gamma\lesssim 1$, the loss rate $\kappa$ becomes larger with increasing $\gamma$. In contrast, in the strong dissipation regime with $\gamma\gtrsim 1$, the loss rate $\kappa$ decreases with increasing $\gamma$, which can be attributed to the suppression of doubly occupied sites due to strong two-body losses, i.e., the Zeno effect Tomita et al. (2017).
@@
Moreover, the loss rate estimated within the SU(3) dTWA is much better than the mean-field result $\hbar\kappa_{\rm MF}=16z\gamma J^{2}U^{-1}/(4+\gamma^{2})\equiv 4z\hbar\Gamma_{\rm eff}$ from a previous work García-Ripoll et al. (2009), where $z=2d=6$ is the coordination number of the lattice geometry. This mean-field result is obtained by performing a site-decoupling approximation for the dynamical equation of $\langle{\hat{n}}_{j}\rangle$ with mapping to a hard-core boson model from the master equation in Eq. (1). The derivation of this effective model is based on a second-order expansion of the hopping term to adiabatically integrate Fock states, including doubly occupied sites, which is validated under the conditions of strong dissipation and strong interaction, i.e., $J\ll\hbar\Gamma_{\rm PA}$ and $J\ll U$. As shown in Fig. 2(c), the mean-field treatment also captures the Zeno effect, and the curve of $\kappa_{\rm MF}$ exhibits a maximum at $\gamma=2$, which is comparable to the dTWA result as well as the experimental result. In particular, the dTWA result and the mean-field result are in good agreement for $\gamma\lesssim 1$. However, the mean-field result overestimates the loss rate for $\gamma\gg 1$. The significant deviations in the strong dissipative regime imply that the mean-field approximation for the hard-core boson representation ignores important effects that suppress loss events more effectively, which are successfully captured in the dTWA.
Figure 3: Rescaled time evolution of the atomic density in units of the effective frequency $\Gamma_{\rm eff}$. (a) The SU(3) dTWA results for $\gamma=1$. The blue solid, red dotted, and black dashed lines correspond to $J/U=0.04$, $0.08$, and $0.1$, respectively. Each line is calculated until $t=100\hbar/U$. For comparison, the mean-field result $n_{\rm MF}(t)=1/(1+4z\Gamma_{\rm eff}t)$ is also plotted by a black solid line. (b) Same as (a) but for $\gamma=5$. (c) Same as (a) but for the SU(3) FP-TWA results. (d) Same as (c) but for $\gamma=5$.
@@
Figure 4: (a) Power-law fitting (green line) of the long-time behavior of the atomic density $n_{j}(t)$ obtained from the SU(3) dTWA (blue circles) for $J=0.08U$ and $\gamma=1$. For comparison, the mean-field result is also shown by the dashed line. The power-law function used for fitting is $f_{\rm power}(t)=b(tU/\hbar)^{-c}$, where $b$ and $c$ are fitting parameters. Note that both the vertical and horizontal axes are in logarithmic scale. (b) Same as (a), but for $\gamma=5$. (c) Fitting parameters $b$ and $c$ in the fitting function $f_{\rm power}(t)=b(tU/\hbar)^{-c}$ for $J=0.08U$ with varying values of $\gamma$, obtained from the results evaluated by the SU(3) dTWA (blue circles) and the SU(3) FP-TWA (red circles).
Let us further discuss the qualitative difference between these two semiclassical descriptions by comparing the results in more detail. The onsite linearity of the SU(3) dTWA leads to long-time behavior for $tU/\hbar\gtrsim 10$ that is qualitatively distinguishable from that in the SU(3) FP-TWA. In particular, for the typical values of $\gamma$ in Fig. 2, the density $n_{j}(t)$ at, for example, $tU/\hbar=100$ in the SU(3) dTWA significantly deviates from that in the SU(3) FP-TWA, even though the short-time behavior agrees well for $tU/\hbar\lesssim 10$. Moreover, as shown in Fig. 2(b), $n_{j}(t)$ obtained by the SU(3) FP-TWA for $\gamma=1$ and $5$ crosses at a certain time, while there is no such intersection found for the SU(3) dTWA results in Fig. 2(a). Furthermore, in Figs. 4(a) and 4(b), we observe that the long-time behavior of the density $n_{j}(t)$ predicted by the SU(3) dTWA can be well fitted with a power-law decay function with an exponent $c<1$, implying that $n_{j}(t)\approx b(tU/\hbar)^{-c}$ for $tU/\hbar\gg 1$. This power-law behavior differs from the mean-field result, indicating that the SU(3) dTWA captures nontrivial effects beyond those described by the simple mean-field approach. While the SU(3) FP-TWA results can also be fitted using the same power-law function, the exponents differ significantly, as shown in Fig. 4(c).
To gain insight into the validity of the TWA descriptions, we will compare the semiclassical results for a small cluster with the numerically exact results of the master equation in Sec. IV.4. There, we find that the long-time behavior obtained by the SU(3) dTWA is qualitatively similar to the exact results, and that the SU(3) FP-TWA is less qualitative than the SU(3) dTWA due to the nonlinearity of the dissipative contribution. Therefore, we conclude that the SU(3) dTWA results in Fig. 2 are more reliable.
@@ IV.2 Linear sweep time sequence of the hopping
Figure 5: (a), (b) SU(3) dTWA results of the nonunitary time evolution of the atomic density $n_{j}(t)$ for the atom loss dynamics, started from the unit filling Mott-insulating state $|\psi_{\rm Mott}\rangle$ at time $t=0$, with the linear sweep time sequence crossing the ground-state phase transition point of the Bose-Hubbard Hamiltonian, i.e., $J_{\rm ini}=0$ and $J_{\rm fin}=0.04U$. The values of the dissipation strength $\gamma$ are indicated in the figures. The time grid size for the Runge-Kutta scheme is set to be $\Delta t=2.0\times 10^{-3}\hbar/U$. (c), (d) Same as (a) and (b) but for SU(3) FP-TWA results. Here, the time grid size for the Euler-Maruyama scheme is set to be $\Delta t=2.0\times 10^{-4}\hbar/U$. For comparison, the experimental results in Ref. Tomita et al. (2017) are also plotted by circles with error bars in (a)–(d). (e), (f) Time evolution of the phase coherence extracted from the single-particle correlation functions, i.e., $S({\bm{0}},t)$, for the linear sweep time sequence evaluated by (e) the SU(3) dTWA and (f) the SU(3) FP-TWA.
@@
Next, we analyze how the two-body loss dissipation affects the coherent time evolution caused by a slow parameter change in the Hamiltonian. At zero temperature and for a commensurate mean filling, the Bose-Hubbard Hamiltonian $\hat{H}_{\rm BH}$ exhibits a second-order quantum phase transition between a Mott-insulating state and a superfluid state Fisher et al. (1989); Capogrosso-Sansone et al. (2007). In the three-dimensional setup with the mean occupation tuned to unit filling, the phase transition occurs at $J_{\rm c}=0.0341U$. This exact critical value was numerically estimated in Ref. Capogrosso-Sansone et al. (2007) by employing a quantum Monte Carlo method. Note that the mean-field Gutzwiller approximation yields the critical point $J_{\rm c}^{\rm MFA}=0.0286U$ van Oosten et al. (2001), which is smaller than the exact value due to the less accuracy of the site-decoupling approximation.
@@
Here, in our simulations, we prepare the unit filling Mott-insulating state $|\psi_{\rm Mott}\rangle$ at $t=0$ and then slowly drive it with the Lindblad equation for a time-dependent hopping amplitude at a fixed value of the dimensionless dissipation parameter $\gamma$. To realize a setup that is close to the real experiment, we assume a linear sweep time sequence of the hopping given by $J(t)=J_{\rm ini}+(J_{\rm fin}-J_{\rm ini})(t/\tau)$ for $0\leq t\leq\tau$ Tomita et al. (2017). Here, $J_{\rm ini}$ and $J_{\rm fin}$ are the initial and final hopping amplitudes, which are set in the Mott-insulating and superfluid regimes, respectively. During this sweep, the onsite interaction strength $U$ remains constant. In the experiment reported in Ref. Tomita et al. (2017), the tuned parameter is the optical lattice depth, which is linearly ramped down from the Mott-insulating regime to the superfluid regime. The dominant effect arising in the dynamics during the ramping down is the exponential increase of the hopping energy. Hence, the linear sweep time sequence of the hopping provides a reasonable simplification of the time sequence realized in the experiment.
@@
Figures 5(a) and 5(b) show the SU(3) dTWA results of the atomic density $n_{j}(t)$ for the atom loss dynamics with the time sequence of the hopping amplitude $J(t)$ described above. We specifically choose $J_{\rm ini}=0$ and $J_{\rm fin}=0.04U$ such that $J_{\rm ini}<J_{\rm c}^{\rm MFA}<J_{\rm c}<J_{\rm fin}$ is satisfied. The sweep time is taken to be $\tau=100\hbar/U$, for which non-adiabatic excitations of the system are sufficiently suppressed during the sweep. For the weak dissipation strength such as $\gamma=0.5$, the loss of atoms becomes significant at an onset time $tU/\hbar\sim 83.5$. Here, the onset time is defined as the time when the atomic density drops below $0.8$, as indicated by a dotted line in Fig. 5(a). As shown in Fig. 5(a), the onset time shifts to an earlier time with increasing $\gamma$ for $\gamma\lesssim 2$. Indeed, the onset times for $\gamma=1$ and $\gamma=2$ are $tU/\hbar\sim 73.9$ and $tU/\hbar\sim 71.1$, respectively. However, for the strong dissipation strength, i.e., $\gamma\gtrsim 2$, the onset time shifts to a later time with increasing $\gamma$, as shown in Fig. 5(b). This result manifests the continuous quantum Zeno effect in the sweep dynamics, and a similar behavior has also been observed experimentally in Ref. Tomita et al. (2017). The threshold strength of $\gamma$ separating the Zeno and non-Zeno regimes is $\gamma\sim 2$, which also agrees well quantitatively with the experimental observation Tomita et al. (2017). Since the numerically exact result for a small cluster also supports a similar onset value, as demonstrated in Sec. IV.4, our result here is considered to be qualitatively reliable.
@@
The SU(3) FP-TWA also describes the delay of the atomic loss in the sweep dynamics for the strong dissipation strength. However, it gives rise to a qualitatively different result, as shown in Figs. 5(c) and 5(d). Namely, in contrast to the dTWA result, the delay occurs already for $\gamma\gtrsim 0.1$, implying that the Zeno regime is broadened on the side of smaller $\gamma$. For instance, along the time slice at $t=40\hbar/U$ in Figs. 5(c) and 5(d), the atomic density is found to be significantly suppressed even for $\gamma=0.1$, and it starts to increase with increasing $\gamma$ beyond the threshold. In other words, the nonlinearity of the FP-type approximation manifests itself in an excessive suppression of the atomic density, thus broadening the Zeno regime in the SU(3) FP-TWA.
@@
In Figs. 5(a)–5(d), we also plot the experimental results of the atomic density for $\gamma=0.5$, $1.1$, $2.3$, and $4.2$ reported in Ref. Tomita et al. (2017), which should be directly compared with our TWA results for $\gamma=0.5$, $1$, $2$, and $4$, respectively. Here, the experimental results are plotted along the time axis such that the values of $J/U$ during the sweep in the experiment are the same as the corresponding values in the linear sweep time sequence employed in our simulations. As shown in these figures, we find that, specifically in the SU(3) dTWA, the qualitative behavior of the $\gamma$ dependence of the melting dynamics is reasonably captured within the experimental error bars.
@@
We further analyze the suppression of the superfluid phase coherence in the presence of the strong two-body loss dissipation. For this purpose, we consider a spatial Fourier transform of the equal-time single-particle correlation function, i.e., the momentum distribution, defined by
@@
where ${\bf R}_{i}$ represents the real-space vector specifying site $i$ and ${\bf k}$ denotes the momentum vector in three dimensions. Note that this distribution is measurable in experiments by using the time-of-flight (TOF) imaging techniques Lewenstein et al. (2012). Here, we specifically set ${\bf k}=(0,0,0)$ for the purpose of quantifying the emergence of coherence over long distances after the hopping sweep. In addition, we note that no inhomogeneity effect due to a trapping potential for the gas is taken into account here.
Figure 5(e) shows the SU(3) dTWA results for $S({\bf 0},t)$. In the absence of dissipation, i.e., $\gamma=0$, the growth of long-distance correlations is recovered within this approximation, implying a dynamical phase transition of the system from the Mott-insulating phase to a superfluid phase. For the weak dissipation strength, i.e., $\gamma\leq 1$, this semiclassical method remains to produce large occupation of the zero-momentum distribution at the final state, i.e., $L^{-3}S({\bf 0},t=\tau)\gg 1$. Hence, the phase coherence comprised of coherent bosons survives well against the weak dissipation. However, for the strong dissipation strength, i.e., $\gamma\gtrsim 2$, the distribution is found to be strongly suppressed such that $L^{-3}S({\bf 0},t=\tau)\sim 1$, implying that the strong dissipation stabilizes the Mott-insulating state and leads to the delay of the evolution toward a superfluid phase due to the continuous Zeno effect. This result is consistent with the experimental observation in TOF absorption images Tomita et al. (2017), revealing that the phase coherence associated with superfluidity is suppressed by the strong dissipation.